Jenniferperez2022:

A group of 8 friends (5 girls and 3 boys) plans to watch a movie, but they have only 5 tickets. How many different combinations of 5 friends could possibly receive the tickets?

surjithayer:

8c5=?

surjithayer:

\[=\frac{ 8\times 7\times 6 \times 5\times 4 }{ 5\times4\times3\times2\times1}=?\]